Find the indicated probability. A hotel has 20 rooms on the first floor, 60 rooms on the second floor, 40 rooms on the third floor, and 20 rooms on the fourth floor. If a hotel room is assigned at random, what is the probability of getting a room on an odd floor? 1/2
Answer #1

Total number of rooms = 20+60+40+20 =140

Total number of rooms in odd floors =20+40=60

If a hotel room is assigned at random then the probability of getting a room on an odd floor is 60 3 140 7